### v1

Abstract
chiappori2025changes study several indices that measure the assortativeness of a matching, including the aggregate likelihood ratio. We provide a counterexample showing that their axiomatization of the aggregate likelihood ratio is not valid as stated. We identify the exact class of indices characterized by the axioms in chiappori2025changes. We then show that the axiomatization of the aggregate likelihood ratio can be recovered by adding new axioms.
1 Introduction
The degree of assortative matching between men and women—how similar partners are in terms of education, age, income, and other socioeconomic characteristics—has long been viewed as a central determinant of inequality and welfare in modern economies (chiappori2023mating). If high-income or highly educated individuals increasingly marry one another, household income inequality may rise even when individual wage distributions remain stable. Understanding the extent and evolution of assortative matching is therefore crucial for interpreting long-run trends in inequality and social mobility.
In a pioneering study, chiappori2025changes propose an axiomatic foundation for indices of assortative matching and provide empirical evidence on changes in marital sorting in the United States. Their framework offers a valuable basis for interpreting trends in marital sorting in subsequent work. In this note, we show that the axiomatization of one measure, the aggregate likelihood ratio, is not correct as currently stated.111eika2019educational define a type-specific likelihood ratio as the ratio of the observed probability that a man and a woman of the same type (e.g., age) marry to the corresponding probability under random matching. The aggregate likelihood ratio is the weighted average of these type-specific likelihood ratios, with weights given by the expected mass of like-type couples under random matching. The measure has been applied to sociology and demographic studies such as pesando2021educational. We identify the exact class of indices characterized by the axioms in chiappori2025changes (Theorem 1). We then propose a refinement that corrects the statement and restores the intended result, while preserving the spirit and practical usefulness of their original approach (Theorem 2).
2 Model
Our notation follows that of chiappori2025changes. We consider two-type markets, i.e., there are two types of men and two types of women. Let $\mathcal{M}=\mathbb{R}_{\geq 0}^{4}\backslash\{\mathbf{0}\}$ denote the entire collection of possible matching patterns. We represent a matching pattern $M\in\mathcal{M}$ by a $2\times 2$ matrix given by
[…]
If the two types are “high income” and “low income”, then $a$ represents the number of matches between high-income men and high-income women, $b$ represents the number of matches between high-income men and low-income women, $c$ represents the number of matches between low-income men and high-income women, and $d$ represents the number of matches between low-income men and low-income women.
We call $M\in\mathcal{M}$ a matching matrix, or simply a matching. We often represent a matching in the vector form as $M=(a,b,c,d)$. For each $M\in\mathcal{M}$, let $|M|:=a+b+c+d$ denote the population size. We write $M>\mathbf{0}$ to mean that every entry of $M$ is positive. Such $M$ is called a positive matching matrix. Let $\mathcal{M}_{>0}$ denote the set of all positive matching matrices.
An index is a function $I:\mathcal{M}\rightarrow\mathbb{R}_{\geq 0}$. For each $M\in\mathcal{M}$, $I(M)$ represents how assortative the matching matrix is. Among several indices considered in chiappori2025changes, we focus on the following one.
  • The aggregate likelihood ratio (abbreviated ALR) is defined by
    $I_{L}(M)=\frac{a+d}{|M|}/\left(\frac{a+b}{|M|}\frac{a+c}{|M|}+\frac{d+b}{|M|}\frac{d+c}{|M|}\right)\text{ for all }M\in\mathcal{M}.$
    Equivalently,
    $I_{L}(M)=(a+d)/\left(\frac{(a+b)(a+c)}{|M|}+\frac{(d+b)(d+c)}{|M|}\right)\text{ for all }M\in\mathcal{M}.$
Notice that ALR is not well-defined for matching matrices $M=(a,b,c,d)$ such that $(a+b)(a+c)+(d+b)(d+c)=0$ (this equation holds if $a=b=d=0$ or $a=c=d=0$). Therefore, in the following analysis, we confine our attention to the class of matching matrices $\tilde{\mathcal{M}}\subseteq\mathcal{M}$ for which ALR is well-defined.
$\tilde{\mathcal{M}}=\{(a,b,c,d)\in\mathcal{M}\mid(a+b)(a+c)+(d+b)(d+c)\neq 0\}.$
We revisit the axioms in chiappori2025changes. Consider a domain $\mathcal{M}^{\prime}\subseteq\mathcal{M}$ with $\mathcal{M}^{\prime}\neq\emptyset$. We define axioms for an index $I:\mathcal{M}^{\prime}\rightarrow\mathbb{R}_{\geq 0}$. [^2] We note that chiappori2025changes’s (chiappori2025changes) define axioms for an index $I:\mathcal{M}\rightarrow\mathbb{R}_{\geq 0}$. The following three axioms are called “invariance axioms”.
  • Scale Invariance. For every $M\in\mathcal{M}^{\prime}$ and $\lambda>0$ with $\lambda M\in\mathcal{M}^{\prime}$, we have $I(M)=I(\lambda\cdot M)$.
  • Side Invariance.
[…]
  • Type Invariance.
[…]
Monotonicity axioms state that if the underlying matching matrix changes in a certain way, then the assortativeness increases.
  • Marginal Monotonicity. For every pair of positive matching matrices $M=(a,b,c,d)\in\mathcal{M}^{\prime}$ and $M^{\prime}=\left(a^{\prime},b^{\prime},c^{\prime},d^{\prime}\right)\in\mathcal{M}^{\prime}$ with the same marginal distributions (i.e., $a+c=a^{\prime}+c^{\prime},a+b=a^{\prime}+b^{\prime},d+b=d^{\prime}+b^{\prime}$, and $\left.d+c=d^{\prime}+c^{\prime}\right)$, $I(M)>I(M^{\prime})$ if and only if $a>a^{\prime}$ (equivalently, $b<b^{\prime},c<c^{\prime}$, or $\left.d>d^{\prime}\right)$.
To axiomatize ALR, chiappori2025changes introduced the following axiom.
  • Random Decomposability. For every pair of positive matching matrices $M,M^{\prime}\in\mathcal{M}^{\prime}$,
    $I\left(M+M^{\prime}\right)=\frac{r(M)}{r\left(M+M^{\prime}\right)}I(M)+\frac{r\left(M^{\prime}\right)}{r\left(M+M^{\prime}\right)}I\left(M^{\prime}\right),$
    where, for each $M\in\mathcal{M}^{\prime}$,
    $r(M)$ $:=\frac{a+b}{|M|}\frac{a+c}{|M|}|M|+\frac{d+b}{|M|}\frac{d+c}{|M|}|M|$
    $=\frac{(a+b)(a+c)+(d+b)(d+c)}{|M|}.$ (1)
We reproduce Theorem 3 of chiappori2025changes.
Theorem CDMZ3.
An index satisfies the three invariance axioms, marginal monotonicity, and random decomposability if and only if it is a positive multiple of aggregate likelihood ratio.
3 Counterexample to Theorem CDMZ3
To see that this theorem does not hold for an index $I:\hat{\mathcal{M}}\rightarrow\mathbb{R}_{\geq 0}$, consider the following variant of $I_{L}$.
$I^{\prime}_{L}(M)=\Bigl(a+d+\frac{b}{2}+\frac{c}{2}\Bigr)/\left(\frac{(a+b)(a+c)}{|M|}+\frac{(d+b)(d+c)}{|M|}\right)\text{ for all }M\in\tilde{\mathcal{M}}$.
In the next section we prove that $I^{\prime}_{L}$ satisfies all the axioms in Theorem CDMZ3; see the proof of the “if” direction of Theorem 1. We show that $I^{\prime}_{L}$ is not a positive multiple of $I_{L}$. Suppose, for contradiction, that $I^{\prime}_{L}$ is a positive multiple of $I_{L}$. Then, for every $M,M^{\prime}\in\tilde{\mathcal{M}}$,
$I_{L}(M)>I_{L}(M^{\prime})\Longleftrightarrow I^{\prime}_{L}(M)>I^{\prime}_{L}(M^{\prime}).$ (2)
Consider the following two matrices.
[…]
Then,
$I_{L}(M)=1>\frac{21}{23}=I_{L}\left(M^{\prime}\right),$
$I_{L}^{\prime}(M)=\frac{3}{2}<\frac{35}{23}=I_{L}^{\prime}\left(M^{\prime}\right).$
We obtain a contradiction to (2).
4 Recovering axiomatization of the aggregate likelihood ratio
In this section, we recover the axiomatization of ALR in chiappori2025changes’s (chiappori2025changes). To this end, we first identify the exact class of indices that satisfy the axioms in Theorem CDMZ3.
Theorem 1.
An index $I:\tilde{\mathcal{M}}\rightarrow\mathbb{R}_{\geq 0}$ satisfies the three invariance axioms, marginal monotonicity, and random decomposability if and only if the following two conditions hold:
  • (i)
    On the class of positive matching matrices $\mathcal{M}_{>0}$, there exist $\alpha,\beta\in\mathbb{R}$ with $\alpha>\beta\geq 0$ such that,
    $I\Bigl(\begin{tabular}[]{c|c}$a$&$b$\\ \hline\cr$c$&$d$\end{tabular}\Bigr)=\cfrac{\alpha a+\beta b+\beta c+\alpha d}{r(M)}\text{ for all }M=\Bigl(\begin{tabular}[]{c|c}$a$&$b$\\ \hline\cr$c$&$d$\end{tabular}\Bigr)\in\mathcal{M}_{>0}.$ (7)
  • (ii)
    On the class of matching matrices $\tilde{\mathcal{M}}\setminus\mathcal{M}_{>0}$, $I$ is an index that satisfies the three invariance axioms.
Proof.
Proof of the “if” direction: It is clear that $I$ satisfies all the axioms on $\tilde{\mathcal{M}}\setminus\mathcal{M}_{>0}$, because on this domain, marginal monotonicity and random decomposability have no bite (recall that these axioms are defined only for two positive matching matrices in $\mathcal{M}_{>0}$).
We show that $I$ satisfies all the axioms on the domain $\mathcal{M}_{>0}$. Suppose that there exist $\alpha,\beta\in\mathbb{R}$ with $\alpha>\beta\geq 0$ such that $I$ is given as in (7). This index satisfies side invariance because it is symmetric with respect to $b$ and $c$. Similarly, it satisfies type invariance because it is symmetric with respect to $a$ and $d$. Furthermore, it satisfies scale invariance because, after multiplying a matching matrix $M\in\tilde{\mathcal{M}}$ by $\lambda>0$, both the numerator and the denominator of $I(M)$ are multiplied by $\lambda$. In the following we show that $I$ satisfies marginal monotonicity and random decomposability.
Proof that $I$ satisfies marginal monotonicity: Take $\epsilon\in\mathbb{R}$ and consider two positive matching matrices $M,M_{\epsilon}\in\mathcal{M}_{>0}$ with the same marginal distributions, i.e.,
[…]
Since
$\left|M_{\varepsilon}\right|=(a+\varepsilon)+(b-\varepsilon)+(c-\varepsilon)+(d+\varepsilon)=|M|,$
the population size is the same between the two matching matrices. The following equations also hold:
$\begin{array}[]{ll}(a+\varepsilon)+(b-\varepsilon)=a+b,&(a+\varepsilon)+(c-\varepsilon)=a+c,\\ (d+\varepsilon)+(b-\varepsilon)=d+b,&(d+\varepsilon)+(c-\varepsilon)=d+c.\end{array}$
Therefore, $r(M)=r(M_{\epsilon})$. Moreover,
Numerator of $I(M_{\epsilon})$ $=\alpha\cdot(a+\epsilon)+\alpha\cdot(d+\epsilon)+\beta\cdot(b-\epsilon)+\beta\cdot(c-\epsilon)$
$=\alpha\cdot a+\alpha\cdot d+\beta\cdot b+\beta\cdot c+2\epsilon\cdot(\alpha-\beta)$
Numerator of $I(M)$ $=\alpha\cdot a+\alpha\cdot d+\beta\cdot b+\beta\cdot c.$
Under $\alpha>\beta$, we have
$I(M_{\epsilon})>I(M)\>\Longleftrightarrow\>\epsilon>0,$
as desired.
Proof that $I$ satisfies random decomposability: For each $M=(a,b,c,d)\in\mathcal{M}_{>0}$, we define
$A(M):=\alpha\cdot a+\alpha\cdot d+\beta\cdot b+\beta\cdot c.$
Then
$I(M)=\frac{A(M)}{r(M)}.$
Consider $M,M^{\prime}\in\mathcal{M}_{>0}$ and let $M^{\prime\prime}:=M+M^{\prime}$. Since $A(\cdot)$ is linear in $(a,b,c,d)$,
$A(M^{\prime\prime})=A(M)+A\left(M^{\prime}\right).$
Then,
$\frac{r(M)}{r(M^{\prime\prime})}I(M)+\frac{r\left(M^{\prime}\right)}{r(M^{\prime\prime})}I\left(M^{\prime}\right)=\frac{r(M)}{r(M^{\prime\prime})}\frac{A(M)}{r(M)}+\frac{r\left(M^{\prime}\right)}{r(M^{\prime\prime})}\frac{A\left(M^{\prime}\right)}{r\left(M^{\prime}\right)}=\frac{A(M)+A\left(M^{\prime}\right)}{r(M^{\prime\prime})}=\frac{A(M^{\prime\prime})}{r(M^{\prime\prime})}=I(M^{\prime\prime}),$
which establishes random decomposability.
Proof of the “only if” direction: Let $I$ be an index that satisfies the three invariance axioms, marginal monotonicity, and random decomposability. It is clear that, on the domain of matching matrices $\tilde{\mathcal{M}}\setminus\mathcal{M}_{>0}$, $I$ is an index that satisfies the three invariance axioms.
It suffices to prove that, on the domain $\mathcal{M}_{>0}$, $I$ is given as stated. We define $\tilde{I}:\mathcal{M}_{>0}\rightarrow\mathbb{R}_{\geq 0}$ by
$\tilde{I}(M)=r(M)\cdot I(M)\text{ for all }M\in\mathcal{M}_{>0}.$ (8)
For each $M\in\mathcal{M}_{>0}$, since $r(M)\geq 0$ and $I(M)\geq 0$, we have $\tilde{I}(M)\geq 0$.
For every pair of positive matching matrices $M,M^{\prime}\in\mathcal{M}_{>0}$, it holds that
$\tilde{I}(M+M^{\prime})$ $=r(M+M^{\prime})\cdot I(M+M^{\prime})$
$=r(M)\cdot I(M)+r(M^{\prime})\cdot I(M^{\prime})$
$=\tilde{I}(M)+\tilde{I}(M^{\prime}),$
where the second equality follows from random decomposability. Therefore, $\tilde{I}$ is additive.
For $M\in\mathcal{M}_{>0}$ and $\lambda>0$,
$\tilde{I}(\lambda M)$ $=r(\lambda M)\cdot I(\lambda M)$
$=\lambda\cdot r(M)\cdot I(\lambda M)$
$=\lambda\cdot r(M)\cdot I(M)$
$=\lambda\cdot\tilde{I}(M),$
where the second equality follows from the definition of $r(M)$ (recall (1)) and the third equality follows from scale invariance. Therefore, $\tilde{I}$ satisfies positive homogeneity of degree 1.
Let
[…]
Also let
$\alpha^{\prime}:=\tilde{I}(M_{1}),\>\>\beta^{\prime}:=\tilde{I}(M_{2}),\>\>\gamma^{\prime}:=\tilde{I}(M_{3}),\>\>\delta^{\prime}:=\tilde{I}(M_{4}).$ (9)
By side invariance and type invariance,
$\beta^{\prime}=\gamma^{\prime},\>\>\alpha^{\prime}=\delta^{\prime}.$ (10)
Consider $M\in\mathcal{M}_{>0}$. Since $M_{1},M_{2},M_{3},M_{4}$ form a basis of $\mathbb{R}^{4}$, there exist coefficients $s_{1},s_{2},s_{3},s_{4}$ such that
$M=s_{1}M_{1}+s_{2}M_{2}+s_{3}M_{3}+s_{4}M_{4}.$
The coefficients are explicitly given by
$s_{1}=a-\frac{a+b+c+d}{5},$
$s_{2}=b-\frac{a+b+c+d}{5},$
$s_{3}=c-\frac{a+b+c+d}{5},$
$s_{4}=d-\frac{a+b+c+d}{5}.$ (11)
Define
$t_{i}^{\prime}:=\max\left\{s_{i},0\right\},\quad t_{i}:=\max\left\{-s_{i},0\right\}\quad(i=1,2,3,4)$
For each $i=1,\dots,4$, we have $t_{i},t_{i}^{\prime}\geq 0$ and $t_{i}^{\prime}-t_{i}=s_{i}$. By the latter equation,
$M=\sum_{i=1}^{4}\left(t_{i}^{\prime}-t_{i}\right)M_{i}\quad\Longleftrightarrow\quad M+\sum_{i=1}^{4}t_{i}M_{i}=\sum_{i=1}^{4}t_{i}^{\prime}M_{i}.$ (12)
Since $\tilde{I}$ satisfies additivity and positive homogeneity of degree 1, [^3] We note that the equation below holds including the case that $t_{i}=0$ for some $i=1,\dots,4$.
$\tilde{I}\Bigl(M+\sum_{i=1}^{4}t_{i}M_{i}\Bigr)=\tilde{I}(M)+\sum_{i=1}^{4}t_{i}\tilde{I}(M_{i}).$
Similarly,
$\tilde{I}\Bigl(\sum_{i=1}^{4}t_{i}^{\prime}M_{i}\Bigr)=\sum_{i=1}^{4}t^{\prime}_{i}\tilde{I}(M_{i}).$
The above two displayed equations together with (12) imply
$\tilde{I}(M)=\sum_{i=1}^{4}t^{\prime}_{i}\tilde{I}(M_{i})-\sum_{i=1}^{4}t_{i}\tilde{I}(M_{i})=\sum_{i=1}^{4}s_{i}\tilde{I}(M_{i}).$
By substituting (9), (10), and (11) into the above equation,
$\tilde{I}(M)=\frac{3\alpha^{\prime}-2\beta^{\prime}}{5}a+\frac{-2\alpha^{\prime}+3\beta^{\prime}}{5}b+\frac{-2\alpha^{\prime}+3\beta^{\prime}}{5}c+\frac{3\alpha^{\prime}-2\beta^{\prime}}{5}d.$
By setting
$\alpha:=\frac{3\alpha^{\prime}-2\beta^{\prime}}{5},\;\>\beta:=\frac{-2\alpha^{\prime}+3\beta^{\prime}}{5},$
we obtain
$\tilde{I}(M)=\alpha a+\beta b+\beta c+\alpha d.$
Since $\tilde{I}(M)=r(M)\cdot I(M)$, $I$ is given in the form of (7).
As noted after (8), $\tilde{I}(M)\geq 0$ for all $M\in\tilde{\mathcal{M}}$, which implies $\alpha\geq 0$ and $\beta\geq 0$ hold. Indeed, if $\alpha<0$, by letting $a,d\rightarrow+\infty$ and $b,c\rightarrow 0$, we obtain a matching matrix $M=(a,b,c,d)\in\mathcal{M}_{>0}$ for which $\tilde{I}(M)<0$, a contradiction to $\tilde{I}(M)\geq 0$. Similarly, $\beta<0$ yields a contradiction. It remains to prove that $\alpha>\beta$. Take $\epsilon\in\mathbb{R}$ with $\epsilon>0$. Consider two positive matching matrices $M,M_{\epsilon}\in\mathcal{M}_{>0}$ such that
[…]
As shown in the proof of the “if” part, we have $r(M)=r(M_{\epsilon})$. By marginal monotonicity of $I$,
$\alpha\cdot(a+\epsilon)+\alpha\cdot(d+\epsilon)+\beta\cdot(b-\epsilon)+\beta\cdot(c-\epsilon)>\alpha\cdot a+\alpha\cdot d+\beta\cdot b+\beta\cdot c.$
This inequality implies $2\epsilon\cdot(\alpha-\beta)>0$. Since $\epsilon>0$, we obtain $\alpha>\beta$, as desired. ∎
Theorem 1 suggests that, to pin down the class of positive multiples of ALR, we need additional axioms that force $\beta=0$ in (7). We consider an axiom that is interpreted as the opposite of the maximum homogamy axiom in chiappori2025changes (defined in Section 5). As in Section 2, we define axioms for an index $I:\mathcal{M}^{\prime}\rightarrow\mathbb{R}_{\geq 0}$, where $\mathcal{M}^{\prime}\subseteq\mathcal{M}$, $\mathcal{M}^{\prime}\neq\emptyset$.
  • Maximum Heterogamy. For each $M\in\mathcal{M}^{\prime}$ and $(a,b,c,d)\in\mathcal{M}^{\prime}$ such that $b>0,c>0,a=d=0$, we have $I(M)\geq I(a,b,c,d)$.
We also introduce the continuity axiom, which states that if the underlying matching matrix changes slightly, then the index changes slightly as well.
  • Continuity: $I:\mathcal{M}^{\prime}\rightarrow\mathbb{R}_{\geq 0}$ is a continuous function.
Theorem 2.
An index $I:\tilde{\mathcal{M}}\rightarrow\mathbb{R}_{\geq 0}$ satisfies the three invariance axioms, marginal monotonicity, random decomposability, maximum heterogamy, and continuity if and only if there exists $\alpha\in\mathbb{R}$ with $\alpha>0$ such that
[…]
Proof.
Proof of the “if” direction: Suppose that there exists $\alpha>0$ such that $I$ is given as in (17). By following the same argument as in the “if” direction of Theorem 1, we can show that $I$ satisfies the three invariance axioms, marginal monotonicity, and random decomposability. Since $I$ is a linear function, it satisfies continuity. For each $(a,b,c,d)\in\tilde{\mathcal{M}}$ with $b>0,c>0,a=d=0$, we have $I(a,b,c,d)=0$. Therefore, maximum heterogamy holds.
Proof of the “only if” direction: By Theorem 1, there exist $\alpha,\beta\in\mathbb{R}$ with $\alpha>\beta\geq 0$ such that $I$ is given as in (7) on the domain $\mathcal{M}_{>0}$. If $I$ satisfies maximum heterogamy, then $\beta=0$. Indeed, if $\beta>0$, by letting $b,c\rightarrow+\infty$ while keeping $a=d=0$, the value of $I(a,b,c,d)$ grows arbitrarily large, contradicting maximum heterogamy.
Note that for any $M\in\tilde{\mathcal{M}}\setminus\mathcal{M}_{>0}$, there exists a sequence $\{M^{k}\}_{k=1}^{\infty}\subseteq\mathcal{M}_{>0}$ such that $M^{k}\rightarrow M$. Therefore, by continuity, the value of $I(M)$ for $M\in\tilde{\mathcal{M}}\setminus\mathcal{M}_{>0}$ is also given by (17). ∎
5 Concluding remarks
We have shown that Theorem 3 of chiappori2025changes does not hold as currently stated and we proposed a modification. We note that Theorem 2 of chiappori2025changes does not hold as currently stated either. Specifically, chiappori2025changes consider the following index called the normalized trace:
$I_{tr}(M)=\left\{\begin{array}[]{cl}1&\text{ if }bc=0,\\ \frac{a+d}{a+b+c+d}\in(0,1)&\text{ if }abcd\neq 0,\\ 0&\text{ if }ad=0.\end{array}\right$.
We remark that this index is not well-defined on the entire domain of matchings $\mathcal{M}$; the conditions “bc=0bc=0” (first line) and “ad=0ad=0” (third line) overlap, so a matching matrix such as $(1,1,0,0)$ satisfies both clauses and is assigned the values 1 and 0 simultaneously. Therefore, we restrict the domain of an index to the set of matchings $\hat{\mathcal{M}}$ defined by
$\hat{\mathcal{M}}=\{M\in\mathcal{M}\mid bc\neq 0\text{ or }ad\neq 0\}.$
The following is Theorem 2 of chiappori2025changes.
Theorem CDMZ2.
The normalized trace is the unique index, up to positive affine transformation, that satisfies the three invariance axioms, marginal monotonicity, maximum homogamy, and population decomposability.
The last two axioms in the statement are defined as follows:
  • Maximum Homogamy. For each $M\in\hat{\mathcal{M}}$ and $(a,b,c,d)\in\hat{\mathcal{M}}$ such that $ad>0$ and $bc=0$, we have $I(a,b,c,d)\geq I(M)$.
  • Population Decomposability. For each pair of positive matchings $M,M^{\prime}\in\hat{\mathcal{M}}$,
    $I\left(M+M^{\prime}\right)=\frac{|M|}{|M|+\left|M^{\prime}\right|}I(M)+\frac{\left|M^{\prime}\right|}{|M|+\left|M^{\prime}\right|}I\left(M^{\prime}\right).$
To see that Theorem CDMZ2 does not hold, consider the following index defined on $\hat{\mathcal{M}}$.
$I^{\prime}_{tr}(M)=\left\{\begin{array}[]{cl}1&\text{ if }bc=0,\\ \frac{a+d+\frac{b}{2}+\frac{c}{2}}{a+b+c+d}\in(0,1)&\text{ if }abcd\neq 0,\\ 0&\text{ if }ad=0.\end{array}\right$.
One can verify that this modified index satisfies the three invariance axioms, marginal monotonicity, maximum homogamy, and population decomposability. [^4] This modified index also satisfies other two monotonicity axioms in chiappori2025changes, namely diagonal monotonicity and off-diagonal monotonicity. However, this modified index is not a positive affine transformation of the normalized trace. To see this, let $M=(a,b,c,d)\in\hat{\mathcal{M}}$. We consider two cases.
Case 1: Suppose $bc=0,ad>0$. Then,
$I_{tr}(M)=1,\quad I_{tr}^{\prime}(M)=1$
Hence
$1=\alpha\cdot 1+\beta\quad\Rightarrow\quad\alpha+\beta=1.$
Case 2: Suppose $ad=0,bc>0$. Then
$I_{tr}(M)=0,\quad I_{tr}^{\prime}(M)=0$
Hence
$0=\alpha\cdot 0+\beta\quad\Rightarrow\quad\beta=0.$
Together with $\alpha+\beta=1$, we have $\alpha=1$. This means that
$I_{tr}(M)=I^{\prime}_{tr}(M)\text{ for all }M\in\mathcal{M}.$ (18)
Let $M=(1,1,1,1)$. Then,
$I_{tr}(M)$ $=\frac{1}{2},$
$I_{tr}^{\prime}(M)$ $=\frac{3}{4}$
We obtain a contradiction to (18).
By adopting a technique similar to that developed in Section 4, we can recover the axiomatization of the normalized trace. We will provide details in an updated version of this note.